If any principal amount is double after 12 years at simple interest then after how much time this principal amount will be three fold

Grade:11

1 Answers

Arun
6 years ago
Let the principal be P and the rate of interest be R% p.a. Then, in time = 12 years, Simple Interest = Principal = P.
Therefore, P = P x R x 12 / 100 => R = (100 / 12)% = (25/3)%
To obtain 3 times the principal, interest = 3P - P = 2P.
Thus, 2P = P x (25 / 3) x T / 100
=> T = 24
Thus, it will take 24 years to get 3 times the invested amount.